OpenStudy (anonymous):

how much money has to be invested at 2.3% interest compounded continuously to have $41,000 after 17 years

terenzreignz (terenzreignz):

How were you taught to do this?

OpenStudy (anonymous):

im not sure how to do this.

terenzreignz (terenzreignz):

Very well ^^ When interest is compounded *continuously*, it follows this formula: Suppose we start with the principal (or initial) amount, P. And we have the interest rate, r, compounded continuously for a number of years, t. Then the final amount, A, is given by the following formula: \[\huge A = P e^{rt}\] Where e is the constant approximately 2.71828

OpenStudy (anonymous):

What is the answer?

terenzreignz (terenzreignz):

Why don't we find out? Here, we're asked for the amount to start with, that is, the principal, so P is what we're looking for. What is the final amount, A?

OpenStudy (anonymous):

im not sure because i do not get the formula at all. please help.

terenzreignz (terenzreignz):

I'm explaining the formula now ^^ My question right now has nothing to do with the formula, my question is: What is the amount we want to end up with?

OpenStudy (anonymous):

$41,000 after 17 years .. Here are the answer choices.. A. $27,731.59 B. $27,741.97 C. $27,762.66 D. $27,793.53

terenzreignz (terenzreignz):

Never mind the answer choices for now. So the amount we want to end up with, A, this is 41,000 And the time is 17 years. So let's substitute accordingly... \[\huge \color{blue}{41000} = P e^{r(\color{red}{17})}\] Now what about the continuously compounded rate of interest, r? What is it?

OpenStudy (anonymous):

2.3%

terenzreignz (terenzreignz):

Which is 0.023 \[\huge {41000} = P e^{\color{green}{0.023}({17})}\]

terenzreignz (terenzreignz):

Now, I hope you have your calculator handy, because you'll need it... How do you get P alone on one side? As in... how do you get rid of everything else but P on the right-side of the equation?

OpenStudy (anonymous):

I DONT UNDERSTAND THIS AT ALL

terenzreignz (terenzreignz):

No need to shout ^^ Suppose we have a simple \[\Large x = yz\] If we wanted to get the y alone on the right-side of the equation, all we would need to do is divide both sides of the equation by z, giving us... \[\Large \frac x z = \frac {yz}z\]\[\Large \frac xz = \frac{y\cancel{z}}{\cancel{z}}\]\[\Large \frac xz = y\]

terenzreignz (terenzreignz):

\[\huge {41000} = P e^{{0.023}({17})}\] Similarly, here, all we need to do is divide both sides of the equation by the bit on the right side that ISN'T P. That is to say, \(\Large e^{0.023(17)}\) Thus giving us: \[\huge \frac{41000}{e^{0.023(17)}} = \frac{P e^{{0.023}({17})}}{e^{0.023(17)}}\] \[\huge \frac{41000}{e^{0.023(17)}} = \frac{P \cancel{e^{{0.023}({17})}}}{\cancel{e^{0.023(17)}}}\] \[\huge \frac{41000}{e^{0.023(17)}} = P\] That wasn't so hard, was it? ^^

OpenStudy (anonymous):

I still dont understand because none of the answer choices match the answer on my calculator..

OpenStudy (anonymous):

Nevermind. Thank you so much <3 !

OpenStudy (anonymous):

You were very helpful :)